OpenStudy (anonymous):

The formulas Cgas=12,000+700x Csolar+30,000+150x model the total cost, in dollar, for gas and solar heating systems x years after installation. What is the ratio of the total cost for gas heating to the total cost for solar heating 5 year after installation? What is this ratio 40 years after installation?

OpenStudy (agreene):

would you like ratios in A:B or \[\frac{A}{B}\] form?

OpenStudy (anonymous):

Yes, please agreena.

OpenStudy (agreene):

I guess it doesnt really matter, lol. for the first case sub in 5 for x and evaluate: 12,000+700(5) = A = 15500 30,000+150(5) = B = 30750 Ratios then being 15500:30750 or \[\frac{15500}{30750}=\frac{62}{123}\] For the second set of ratios, plug in x=40 and evaluate.

OpenStudy (anonymous):

Thank you very much agreena. :)